# Bayesian Estimation and Tuning-Free Rank Detection for Probability Mass Function Tensors

### v1

@@ IV-B Computing the natural gradient of the ELBO
4:     Randomly sample data points $\{\bm{y}_{t_{m}}\}_{m=1}^{M}$ from $\bm{Y}$.
5:     Update $\rho_{r,t_{m}},\,\forall r,t_{m}$ using (32), (34), and (36).
6:     Compute noisy gradients ^$\widehat{\nabla}_{\widetilde{\alpha}_{\lambda,r}}\mathcal{L}_{m}$ and
       ^$\widehat{\nabla}_{\widetilde{\alpha}_{n,r,i_{n}}}\mathcal{L}_{m}$, $\forall m$ using (50) and (51), respectively.
7:     Compute the learning rates $\xi_{\lambda}$ and $\xi_{n,r},\forall n,r$
       (e.g., (54)).
8:     Update $\widetilde{\alpha}_{\lambda,r},\,\forall r$ using (52).
9:     Update $\widetilde{\alpha}_{n,r,i_{n}},\,\forall n,r,i_{n}$ using (53).
10: until the log-likelihood of a held-out dataset converges.
11: Find $\widehat{\bm{\lambda}}$ and $\{\widehat{\bm{A}}_{n}\}_{n=1}^{N}$ using (24) and (30), respectively.
12: Find $\widehat{R}$ by deleting components corresponding to
$\quad\widehat{\lambda}_{r}<\epsilon$.
IV-C Stochastic optimization of the ELBO
The natural gradients in (48) and (49) can be used directly in a steepest ascent algorithm to maximize the ELBO. However, this approach would have the same computational complexity as the VB-PMF algorithm since it would require analyzing the entire dataset and computing the local variational parameters for each data point. This problem is solved using stochastic optimization, a technique which follows noisy estimates of the gradient in order to optimize an objective function. The noisy gradients are cheaper to compute since they are estimated by subsampling the data.
A noisy natural gradient is constructed by sampling $M\ll T$ data points at random, computing the natural gradient for each data point as though it was replicated $T$ times, and finally averaging over all $M$ data points in the minibatch. It was shown in [22] that this approach yields an unbiased estimate of the natural gradient. Let the random data points be $\{\bm{y}_{t_{1}},\dotsc,\bm{y}_{t_{M}}\}$, where $t_{m}\sim U(1,T)$ and $\bm{y}_{t_{m}}=[y_{1,t_{m}},\dotsc,y_{N,t_{m}}]^{\mathsf{T}}$ for $m=1,\dotsc,M$. Thus, proceeding from (48) and (49), the noisy natural gradients of the ELBO as a function of the global variational parameters $\widetilde{\bm{\alpha}}_{\lambda}$ and $\widetilde{\bm{\alpha}}_{n,r}$ at the $m$-th data point and the $k$-th iteration are given by
^$\widehat{\nabla}_{\widetilde{\alpha}^{(k)}_{\lambda,r}}\mathcal{L}_{m}=\alpha_{\lambda}+T\rho^{(k)}_{r,t_{m}}-\widetilde{\alpha}^{(k-1)}_{\lambda,r},$ (50)
^$\widehat{\nabla}_{\widetilde{\alpha}^{(k)}_{n,r,i_{n}}}\mathcal{L}_{m}=\alpha_{n,r}+I\{y_{n,t_{m}}=i_{n}\}T\rho^{(k)}_{r,t_{m}}-\widetilde{\alpha}^{(k-1)}_{n,r,i_{n}},$ (51)
where the indicator function $I\{\cdot\}$ yields 1 when $y_{n,t_{m}}=i_{n}$ and 0 otherwise.
These noisy gradients can now be used in a steepest ascent algorithm. For the $k$-th iteration, we have the following updates
$\widetilde{\alpha}_{\lambda,r}^{(k)}=\widetilde{\alpha}_{\lambda,r}^{(k-1)}+\xi^{(k)}_{\lambda}\Bigg{(}\frac{1}{M}\sum_{m=1}^{M}\widehat{\nabla}_{\widetilde{\alpha}_{\lambda,r}^{(k)}}\mathcal{L}_{m}\Bigg{)},$ (52)
$\widetilde{\alpha}_{n,r,i_{n}}^{(k)}=\widetilde{\alpha}_{n,r,i_{n}}^{(k-1)}+\xi^{(k)}_{n,r}\Bigg{(}\frac{1}{M}\sum_{m=1}^{M}\widehat{\nabla}_{\widetilde{\alpha}_{n,r,i_{n}}^{(k)}}\mathcal{L}_{m}\Bigg{)},$ (53)
Figure 2: Joint PMF estimation and rank detection for different number of observations $T$. The true rank of the PMF tensor is $R=5$. (a) Mean estimated rank ($p=0$); (b) Mean estimated rank ($p=0.3$); (c) Mean KLD between the estimated and true PMF tensors ($p=0.3$); (d) Mean runtime ($p=0.3$).
where $\xi^{(k)}_{\lambda}$ and $\xi^{(k)}_{n,r}$ are the learning rates for $\widetilde{\bm{\alpha}}_{\lambda}$ and $\widetilde{\bm{\alpha}}_{n,r}$, respectively. To guarantee convergence, both learning rate sequences should be chosen to satisfy the Robbins-Monro conditions [29], i.e., $\sum_{k}\xi^{(k)}=\infty$ and $\sum_{k}\xi^{(k)^{2}}<\infty$. Alternatively, the authors in [30] have proposed a method to compute the learning rate adaptively, based on minimizing the expected squared norm of the error between the stochastic updates (e.g., (52)) and the optimal variational parameters (e.g., (22)). The learning rate requires no hand-tuning and performs better than the former approach. Therefore, we adopt the method in [30] to compute the learning rates. As an example, the learning rate for the update of $\widetilde{\bm{\alpha}}_{\lambda}$ is computed as follows (a similar procedure holds for $\widetilde{\bm{\alpha}}_{n,r}$). Let $\bm{g}^{(k)}\in\mathbb{R}^{R}$ contain the noisy natural gradients defined in (50) for all $r$. Then, the learning rate at the $k$-th iteration is given by [30]
$\xi_{\lambda}^{(k)}=\frac{\mathbb{E}[\bm{g}^{(k)}]^{\mathsf{T}}\cdot\mathbb{E}[\bm{g}^{(k)}]}{\mathbb{E}[\bm{g}^{(k)^{\mathsf{T}}}\bm{g}^{(k)}]}\approx\frac{\overline{\bm{g}}^{(k)^{\mathsf{T}}}\overline{\bm{g}}^{(k)}}{\overline{\bm{h}}^{(k)}},$ (54)
where the expected values $\mathbb{E}[\bm{g}^{(k)}]$ and $\mathbb{E}[\bm{g}^{(k)^{\mathsf{T}}}\bm{g}^{(k)}]$ are approximated by their respective exponential moving averages $\overline{\bm{g}}^{(k)}$ and $\overline{\bm{h}}^{(k)}$.
The complete algorithm, which we call stochastic VB-PMF (SVB-PMF), is summarized in Algorithm 2. The algorithm has a number of attractive properties. Firstly, all the quantities that need to be computed in SVB-PMF are already in place once VB-PMF is implemented, thus VB-PMF can be converted into its stochastic counterpart without much effort. Secondly, as pointed out earlier, the key advantage of the stochastic approach is scalability as the size of the dataset increases. The overall complexity of the algorithm is $\mathcal{O}(NRM)$, where $M$ is the batch size. Thus, if the batch size is chosen such that $M$ has a sub-linear dependence on $T$ (e.g., $M=T^{1/b},b>1$), considerable savings in computation time can be achieved.
V Results
Figure 3: Joint PMF estimation and rank detection for different number of observations $T$. The true rank of the PMF tensor is $R=10$. (a) Mean estimated rank ($p=0$); (b) Mean estimated rank ($p=0.3$); (c) Mean KLD between the estimated and true PMF tensors ($p=0.3$); (d) Mean runtime ($p=0.3$).
In this section, we evaluate the performance of the proposed Bayesian estimation algorithms through carefully designed synthetic data simulations. In addition, both approaches are applied to movie recommendation data to demonstrate their effectiveness in real statistical learning tasks.
V-A Synthetic-Data Experiments
V-A1 VB-PMF performance
@@
We consider a small-scale experiment with $N=5$ discrete random variables each taking $I_{n}=10$ states. We assume that the 5-dimensional joint PMF tensor $\bm{\mathcal{X}}$ admits a naïve Bayes model. The number of states that the hidden variable $H$ can take is set to $R=\{5,10\}$. The elements of $\bm{\lambda}\in\mathbb{R}^{R}$ and $\{\bm{A}_{n}\}_{n=1}^{N}\in\mathbb{R}^{I_{n}\times R}$ are drawn randomly from $U(0.3,1)$ and $U(0,1)$, respectively, followed by normalization according to the probability simplex constraints. The tensor $\bm{\mathcal{X}}$ is then constructed according to (2) and is fixed for the rest of the experiment. For each trial, a dataset $\bm{Y}\in\mathbb{R}^{N\times T}$ containing observations $\bm{y}_{t}$ is obtained by first sampling $\bm{x}_{t}$ from $\bm{\mathcal{X}}$ (according to the naïve Bayes model), then randomly and independently zeroing out its elements according to the outage probability $p$ (cf. Section II). The values of the outage probability are set to $p=\{0,0.3\}$. The accuracy of the estimate is evaluated in terms of the KLD
Figure 4: VB-PMF rank estimation performance under various values of the hyperparameter $\alpha_{\lambda}$ for true ranks $R=5$ (left) and $R=10$ (right). Here, $T=10^{5}$ and the outage probabilities are $p=\{0,0.3\}$.
Figure 5: SVB-PMF rank estimation performance versus the number of observations $T$ under various minibatch sizes $M$. The true rank is $R=25$. Left: Maximum runtime allowed is 10 hours with the convergence threshold set to $10^{-7}$. Right: Maximum runtime allowed is 15 hours with the convergence threshold set to zero.
between the true PMF $\bm{\mathcal{X}}$ and the estimated PMF $\widehat{\bm{\mathcal{X}}}$, computed as
@@
$D(\bm{\mathcal{X}}\,\|\,\widehat{\bm{\mathcal{X}}})=\sum_{i_{1},\dotsc,i_{N}}\bm{\mathcal{X}}(i_{1},\dotsc,i_{N})\log\frac{\bm{\mathcal{X}}(i_{1},\dotsc,i_{N})}{\widehat{\bm{\mathcal{X}}}(i_{1},\dotsc,i_{N})}.$
The performance of VB-PMF is compared with three model selection techniques for naïve Bayes models [15]: the Akaike information criterion (AIC), the Bayesian information criterion (BIC), and the decomposed normalized maximum likelihood (DNML) criterion. These techniques require cross-validation in order to find the best rank. To estimate the CPD components, we use SQUAREM-PMF [12], an accelerated form of the expectation maximization (EM) algorithm [11], and test candidate ranks in the range $\{1,\dotsc,R+5\}$. We initialize VB-PMF with $R=23$ components [^4] For an $N$-dimensional tensor with full-rank factor matrices $\bm{A}_{n}\in\mathbb{R}^{I_{n}\times R}{\color[rgb]{0,0,0},}\,\forall n$, the Kruskal condition $\sum_{n}\mathrm{min}(I_{n},R)\geq 2R+(N-1)$ is sufficient for a unique decomposition [27]. In our present case, the full-rank condition is satisfied and thus the maximum possible rank is $R=23$. and set the hyperparameters $\alpha_{\lambda}$ and $\alpha_{n,r}$ to $10^{-6}$ and $1$, respectively. This hyperparameter setting favors sparse solutions for $\bm{\lambda}$ and all possible solutions for $\{\bm{A}_{n}\}_{n=1}^{N}$, allowing the rank to be estimated as part of the inference. In addition, the pruning threshold is set to $\epsilon=10^{-3}$
Figure 6: Mean negative log-likelihood (NLL) of a held-out test dataset versus the number of observations $T$. Three SVB-PMF variants with different minibatch sizes $M$ are shown.
Figure 7: Complementary cumulative distribution function (CCDF) of the runtime for different number of observations $T$. The vertical dashed lines indicate the mean runtime. Maximum allowed runtime is 10 hours with the convergence threshold set to $10^{-7}$.
while the convergence thresholds for both SQUAREM-PMF and VB-PMF are set to $10^{-7}$.

### v2

@@ III-D Algorithm: VB-PMF
4:   Update $\rho_{r,t},\,\forall r,t$ using (32), (34), and (36).
5:   Update $\widetilde{\alpha}_{\lambda,r},\,\forall r$ using (22).
6:   Update $\widetilde{\alpha}_{n,r,{\color[rgb]{0,0,0}i}},\,\forall n,r,{\color[rgb]{0,0,0}i}$ using (28).
7:   Evaluate the ELBO using (39).
8: until the ELBO converges
9: Find $\widehat{\bm{\lambda}}$ and $\{\widehat{\bm{A}}_{n}\}_{n=1}^{N}$ using (24) and (30), respectively.
10: Find $\widehat{R}$ by selecting components corresponding to
$\quad\widehat{\lambda}_{r}{\color[rgb]{0,0,0}>\alpha_{\lambda}/T}$.
Figure 2: Typical convergence behavior of the VB-PMF algorithm, initialized with $R=23$ components, for a rank-5 PMF tensor. Left: convergence of the evidence lower bound (ELBO) (left y-axis) and evolution of the rank (right y-axis). Right: elements of the estimated loading vector $\widehat{\bm{\lambda}}$ after convergence.
where we have used (22) and (37) together with the fact that $\sum_{t,r}\rho_{r,t}=T$. If the $r$th component is not responsible for explaining the observations, then $\sum_{t}\rho_{r,t}\approx 0$. Therefore, for such a component, if $\alpha_{\lambda}\rightarrow 0$, then $\widehat{\lambda}_{r}\approx 0$ and the entire rank-one component (i.e., $\widehat{\lambda}_{r}$ and $\{\widehat{\bm{a}}_{n,r}\}_{n=1}^{N}$) can be pruned out of the model. The pruning threshold $\alpha_{\lambda}/T$ is obtained from (40) by noticing that when $\sum_{t}\rho_{r,t}\approx 0$, then $\widehat{\lambda}_{r}\approx\alpha_{\lambda}/T$ given that $T\gg R\alpha_{\lambda}$.
The rank can therefore be determined via a single training run (i.e., without cross-validation) in which the algorithm is initialized with $R>R_{\mathrm{true}}$ components (where $R_{\mathrm{true}}$ is the unknown true rank of the PMF tensor), followed by selecting components with a value greater than $\alpha_{\lambda}/T$ after convergence. One strategy for choosing the initial number of components is to set $R$ as the maximum possible rank for which the Kruskal uniqueness conditions for the CPD are satisfied [34], [35]. Alternatively, the initial rank can be set manually for computational efficiency. The algorithm is deemed to have converged when the difference between successive ELBO values is less than a specified threshold.
Fig. 2 illustrates the typical convergence behavior of the VB-PMF algorithm. Here, we consider a rank-5 PMF with $N=5$ and $I_{n}=10,\forall n$. The low-rank components are estimated from a dataset having $T=10^{4}$ observations and outage probability $p=0$. It can be observed that the ELBO increases monotonically. This corresponds to a successive decrease in the KLD between the estimated and true posterior distributions. On the other hand, the inferred rank decreases from its initial value of 23 components [^4] For an $N$-dimensional tensor with factor matrices $\bm{A}_{n}\in\mathbb{R}^{I_{n}\times R}{\color[rgb]{0,0,0},}\,\forall n$, the Kruskal condition $\sum_{n}{\color[rgb]{0,0,0}k_{\bm{A}_{n}}}\geq 2R+(N-1)$ is sufficient for a unique decomposition [35]. Here, $k_{\bm{A}_{n}}$ denotes the Kruskal rank of the matrix $\bm{A}_{n}$, i.e., the largest integer such that every $k_{\bm{A}_{n}}$ columns are linearly independent. Note that if the columns of $\bm{A}_{n}$ are drawn independently from an absolutely continuous distribution, then $k_{\bm{A}_{n}}=\mathrm{min}(I_{n},R)$ almost surely. Thus, for the setup considered, the maximum possible rank is $R=23$. to converge to the true value. The stem plot shows the expected value of the loading vector $\bm{\lambda}$ after convergence, demonstrating that the desired latent components have been successfully detected while the extraneous components are completely suppressed. Having set $\alpha_{\lambda}=10^{-6}$, the inset plot shows that these extraneous components converge to the value $\alpha_{\lambda}/T=10^{-10}$ as expected.
The complexity of the VB-PMF algorithm consists of updating the parameters of the variational distributions and computing the ELBO to check for convergence. The parameters $\rho_{r,t}$ are computed (per iteration) in $\mathcal{O}(NRT)$ while the combined complexity of computing $\widetilde{\bm{\alpha}}_{\lambda}$ and $\widetilde{\bm{\alpha}}_{n,r}$ is $\mathcal{O}(NR)$. In addition, the complexity of computing the ELBO is $\mathcal{O}(NRT)$. Thus, the overall complexity is $\mathcal{O}(NRT)$ per iteration.
IV Synthetic-Data Experiments
Figure 3: Joint PMF estimation performance of VB-PMF, compared to other PMF estimation approaches, versus the number of observations $T$, averaged over 50 trials.
In this section, we evaluate the performance of the proposed PMF estimation approach using carefully designed synthetic data simulations to demonstrate its effectiveness in joint parameter and rank inference.
@@ IV Synthetic-Data Experiments
We consider a small-scale experiment with $N=5$ discrete random variables each taking $I_{n}=10$ states. We assume that the 5-dimensional joint PMF tensor $\bm{\mathcal{X}}$ admits a naïve Bayes model. The elements of $\bm{\lambda}\in\mathbb{R}^{R}$ and $\{\bm{A}_{n}\}_{n=1}^{N}\in\mathbb{R}^{I_{n}\times R}$ are drawn randomly from $U(0.3,1)$ and $U(0,1)$, respectively, followed by normalization according to the probability simplex constraints. Sampling the elements of $\bm{\lambda}$ from $U(0.3,1)$ ensures that the rank-one component associated with the smallest element in $\bm{\lambda}$ contributes significantly to the rank of the PMF tensor. The tensor $\bm{\mathcal{X}}$ is then constructed according to (2) and is fixed for the rest of the experiment. For each trial, a dataset $\bm{Y}\in\mathbb{R}^{N\times T}$ containing observations $\bm{y}_{t}$ is obtained by first sampling $\bm{x}_{t}$ from $\bm{\mathcal{X}}$ (according to the naïve Bayes model), then randomly and independently zeroing out its elements according to the outage probability $p$ (cf. Section II). The accuracy of the estimate is evaluated in terms of the mean KLD and the mean squared relative error (MSRE) between the true PMF $\bm{\mathcal{X}}$ and the estimated PMF $\widehat{\bm{\mathcal{X}}}$, computed as
@@
${\color[rgb]{0,0,0}\mathrm{KLD}_{\mathrm{mean}}}={\color[rgb]{0,0,0}{\color[rgb]{0,0,0}\hat{\mathbb{E}}}}\left[\sum_{i_{1},\dotsc,i_{N}}\bm{\mathcal{X}}(i_{1},\dotsc,i_{N})\log\frac{\bm{\mathcal{X}}(i_{1},\dotsc,i_{N})}{\widehat{\bm{\mathcal{X}}}(i_{1},\dotsc,i_{N})}\right]$
${\color[rgb]{0,0,0}\mathrm{MSRE}={\color[rgb]{0,0,0}\hat{\mathbb{E}}}\left[\frac{\|\bm{\mathcal{X}}-\widehat{\bm{\mathcal{X}}}\|_{\mathsf{F}}^{2}}{\|\bm{\mathcal{X}}\|_{\mathsf{F}}^{2}}\right],}$
where $\hat{\mathbb{E}}$ denotes the empirical average over all trials.
We initialize VB-PMF with 23 components (see footnote 4) and set the hyperparameters $\alpha_{\lambda}$ and $\alpha_{n,r}$ to $10^{-6}$ and $1$, respectively. This hyperparameter setting favors sparse solutions for $\bm{\lambda}$ and all possible solutions for $\{\bm{A}_{n}\}_{n=1}^{N}$, allowing the rank to be estimated as part of the inference.
The performance of VB-PMF is compared with various PMF estimation algorithms and model order selection techniques as described below:
  • SQ-AIC, SQ-BIC, SQ-DNML: The PMF is estimated using SQUAREM-PMF [13] which is an accelerated form of the expectation maximization (EM) algorithm [12]. The best rank is selected from a list of candidates $\{1,\dotsc,R+5\}$ using the AIC [14], the BIC [15], and the DNML [16] likelihood-based criteria.
  • CTF3D-ValErr, CTF5D-ValErr: The PMF is estimated using the coupled tensor factorization (CTF) approach with a KLD loss criterion [7], [36]. We consider joint PMF recovery based on third-order marginals (CTF3D) as well as the full fifth-order PMF (CTF5D). The best rank is selected from a list of candidates $\{1,\dotsc,R+5\}$ by observing validation errors (ValErr) on a small held-out dataset.
  • SQ-Thresh (Low), SQ-Thresh (High): The PMF is estimated using SQUAREM-PMF [13] initialized with 23 components as with VB-PMF. For SQ-Thresh (low), we remove rank-one components whose estimated weight (probability) is less than 10% of the largest element in the estimated loading vector. Therefore, this approach serves as a naïve baseline for model order selection. On the other hand, for SQ-Thresh (High), we assume that the true rank $R$ is known and keep only the largest $R$ components after estimation. In both cases, we renormalize the CPD components to restore the probability simplex constraints.
We present results averaged over 50 independent trials.
IV-A Performance with varying number of observations
Figure 4: Rank estimation performance versus the outage probability $p$.
Figure 5: Mean KLD between the estimated and true PMF tensors versus the outage probability $p$.
Fig. 3 presents results for a rank-$5$ PMF tensor and an outage probability $p=0$, i.e., no missing observations. We observe that the rank inferred by VB-PMF converges to the true rank as the number of observations is increased. This demonstrates empirically that VB-PMF is statistically consistent [^5] A consistent model selection criterion is one which almost surely (i.e., with probability one) selects the candidate model having the correct dimension as the sample size grows infinitely, given that the dimension of the true model is represented among the candidates [37].. We emphasize that the rank is estimated from the data as part of the inference procedure without cross-validation. Both SQ-BIC and SQ-DNML are consistent criteria and therefore detect the true rank as $T$ increases. On the other hand, SQ-AIC and both CTF approaches exhibit a large variance in the inferred rank. In addition, thresholding as applied in SQ-Thresh (Low) is not able to estimate the correct rank, although the error decreases as $T$ grows.